OpenStudy (anonymous):

Someone help please! Which equation below represents a horizontal line? A) y=4 B) x=4 C) y = x + 0 D) y = x -0

Nnesha (nnesha):

horizontal line doesn't have slope

OpenStudy (anonymous):

Could the answer be A ?

OpenStudy (anonymous):

the answer is a because the y value never changes. it is always 4

Nnesha (nnesha):

yes right where y = b no x no slope

OpenStudy (anonymous):

Okay, that's what I thought. Thank you @koook & @Nnesha :)

Nnesha (nnesha):

my pleasuer :)

OpenStudy (anonymous):

Could you help me with one more?

Nnesha (nnesha):

sure

OpenStudy (anonymous):

Which equation below represents a vertical line? A) y=4 B) x=4 C) y = x + 0 D) y = x - 0

Nnesha (nnesha):

so vertical line which on x'axis

OpenStudy (anonymous):

Huh?

OpenStudy (anonymous):

Ohh, so the correct answer would be B ?

Nnesha (nnesha):

yep right

OpenStudy (anonymous):

Okay, thanks!

Nnesha (nnesha):

where x = b